OpenStudy (anonymous):

Two sets of quadratic equations are shown below in various forms. Line 1: x2 − x − 6 (x + 2)(x − 3) (x + 0.5)2 − 6.25 Line 2: x2 − x − 2 (x + 1)(x − 2) (x − 0.5)2 − 2.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (ipwnbunnies):

The factored form looks good. Sorry, I don't know what the form of the last column is. I don't wanna give you the wrong answer or the wrong path.

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Hmm........

OpenStudy (anonymous):

i think its only line one

OpenStudy (31356):

A. Refer to the attached calculations using the Mathematica program. http://assets.openstudy.com/updates/attachments/52409c5ce4b090015417d3ad-robtobey-1379966644672-nitotheburito.pdf

OpenStudy (anonymous):

what do you think @31356

OpenStudy (31356):

Great observations by the way! :D

OpenStudy (31356):

So it's Line 1 :D

OpenStudy (anonymous):

thanks

OpenStudy (31356):

Thx for the medal :D

OpenStudy (31356):

Anytime :D

OpenStudy (anonymous):

Line 2. Why? (x + 1)(x - 2) is the equivalent factoring form of x^2 - x - 2 (x - 0.5)^2 - 2,25 is another equivalent form (x - 0.5)^2 - 2.25 = x^2 - x + 0.25 - 2.25 = x^2 - x - 2

OpenStudy (anonymous):

thu1935 is right, the correct answer is B. I just got it right.